A new valve has been installed on top of the failed BOP. Tonight, they will be closing the valves and evaluating the response. I believe that something like 8000 psi on the gauge, means success. http://www.bp.com/liveassets/bp_int...ponse/STAGING/local_assets/html/Hos_ROV2.html All live feeds http://www.bp.com/sectionbodycopy.do?categoryId=9034366&contentId=7063636 Edit: In fact, they just now repositioned the camera to center the guage. They must be getting ready.